# ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation

### v1

@@ 4.3 Student Models
4.3.1 Performance Comparison with Baseline
Our student models, trained with explicit and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models. As shown in Table 2,the improvements over the baseline models are statistically significant, with p-values less than 0.05 (indicated by ).
@@ 4.3.1 Performance Comparison with Baseline
4.4 ReasoningRank Student Model vs. Baseliness
Our ReasoningRank student model demonstrates strong performance across various datasets, often outperforming the baseline models. For instance, on the DL19 dataset, our student model achieves an NDCG@5 of 75.39%, surpassing DeBERTa (68.53%), RankVicuna (71.12%), and RankZephyr (72.23%). Similarly, on the DL20 dataset, our model attains 72.36%, outperforming DeBERTa (64.18%) and RankVicuna (68.66%).
On some datasets, our model achieves the best performance among all models. For example, on the News dataset, our student model reaches an NDCG@5 of 53.81%, slightly outperforming RankZephyr (53.33%) and significantly better than DeBERTa (50.24%).
However, there are cases where other models perform better. On the Touche dataset, RankVicuna achieves an NDCG@5 of 48.73%, which is higher than our student model’s 36.22%. This may be due to RankVicuna’s design for zero-shot listwise reranking, which could be particularly effective for argument retrieval tasks represented in Touche. Similarly, on the Robust04 dataset, RankZephyr attains a higher NDCG@5 of 60.72% compared to our model’s 53.50%.
These variations suggest that while our model performs well, certain models may have advantages in specific domains or datasets due to their architectures or training strategies. For example, RankVicuna’s strong performance on Touche might be attributed to its effectiveness in handling argumentation and persuasive texts, which are prevalent in that dataset. RankZephyr’s higher score on Robust04 could be due to its ability to leverage zero-shot capabilities effectively in that particular dataset.
@@ 4.4.1 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
The BRIGHT dataset is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance. Table 3 shows the NDCG@5 results on the BRIGHT dataset. In the Earth Science domain, the student model achieves an NDCG@5 of 27.0%, closely approaching the teacher model’s 31.0%. Similarly, in Psychology, the student model attains 26.0%, compared to the teacher’s 28.9% and BM25’s 11.6%.
Domain BM25 Teacher Student GPT-4 COT
Bio. 17.2 24.3 21.0 28.0
Earth. 24.2 31.0 27.0 37.1
Econ. 13.2 19.0 16.5 22.1
Psy. 11.6 28.9 26.0 34.2
Rob. 12.0 16.0 14.5 19.0
Stack 16.0 21.0 18.5 24.2
Sus. 12.4 18.0 16.2 21.9
Leet. 22.1 27.6 25.0 31.4
Pony 8.5 5.9 5.2 6.7
AoPS 6.0 7.6 6.8 9.2
TheoQ. 9.4 18.5 15.3 23.0
TheoT. 3.5 17.0 14.2 21.7
Average 13.1 19.6 17.5 23.2
@@
4.4.2 Quality of Generated Reasons
@@
Dataset Model BLEU ROUGE-L
DL19 Baseline 18.0 32.5
Ours 21.4 36.8
DL20 Baseline 19.5 34.0
Ours 24.0 38.2
Covid Baseline 16.0 30.0
Ours 18.9 34.5
Touche Baseline 17.5 33.0
Ours 20.5 37.1
News Baseline 20.0 35.0
Ours 22.7 39.3
Table 4: Evaluation of reason generation using BLEU, and ROUGE-L scores for baseline and student models.
4.4.3 Ablation Study: Impact of Reasoning Strategies
Results
In Table 5, by removing comparison reasoning results in a performance drop on DL19 from 75.39 to 74.12 (NDCG@5). Similarly, removing explicit reasoning causes a performance decrease on DL20 from 72.36 to 72.25 (NDCG@5).
Analysis
Both reasoning strategies contribute to performance improvements. The ablation study shows that omitting either reasoning component results in a noticeable decline in performance, highlighting their complementary roles in improving ranking effectiveness.
@@ Results
Comparison & explicit 75.39 / 73.85 72.36 / 71.22
w/ explicit reasoning 74.54 / 73.18 70.11 / 70.97
w/ comparison reasoning 74.12 / 72.94 72.25 / 70.43
Basic w/o reasoning 73.23 / 72.12 70.82 / 69.83
Table 5: Ablation study: Impact of explicit and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
4.4.4 Effects of Training Data Size Across Student Models
Figure 4 i shows the performance of small models trained on 100, 1,000, and 2,000 examples. Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it. We observed similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
@@ 5 Discussion
Our experiments demonstrate that the ReasoningRank framework effectively enhances document reranking by incorporating explicit and comparative reasoning. The student models achieve competitive results with significantly reduced computational resources compared to large teacher models. The statistically significant improvements confirm the effectiveness of reasoning strategies in enhancing model performance.
An important aspect of ReasoningRank is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings. This dual capability suggests that the model could be applied to other tasks that require both ranking and reasoning generation.
While our current work focuses on document reranking in information retrieval, the approach of integrating reasoning generation with ranking decisions could be extended to other domains. For example, in recommendation systems, the model could rank products or services while providing explanations for its recommendations, thereby improving user trust. Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could enhance understandabilityJiang et al. (2024).
However, we acknowledge that our paper does not evaluate the model’s generative abilities on tasks beyond document reranking. This is a limitation in assessing the full scope of the model’s applicability. Future work should explore how the model performs in different contexts to verify its generalizabilityMcElfresh et al. (2022); Rahmani et al. (2022).
@@ 6 Conclusion
In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates explicit and comparative reasoning. Our method leverages the power of large language models to generate interpretable reasons, which are then distilled into smaller, more efficient student models. Our experiments demonstrate the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
@@ Appendix D Model Behavior Analysis
In the reasoning generation tasks, the performance of the model varied across explicit reasoning, comparison reasoning, and their combination. Below, we provide an analysis of unexpected behavior based on 2,000 queries.
@@
The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we observed:
  • Explicit reasoning: 63 duplicate passage identifiers (3%).
@@
Missing documents were most frequent in explicit reasoning:
  • Explicit reasoning: 642 missing documents (32% of queries).
@@ Appendix F Prompts for Reasoning
Table 10: Basic prompt, explicit reasoning prompt, and comparison reasoning prompt.
@@ 3. If there is no direct evidence, try to infer from existing evidence and answer one question:
Explicit Reasoning Prompt:
Then, read every passage one-by-one and find the sentence where the method is the direct answer

### v2

@@ 4.2.2 Student models
Impact of Reasoning Strategies
we also evaluated different reasoning strategies in the R2R student models, specifically analyzing the "w/o Reasoning" (without reasoning), "w/ direct relevance reasoning" (with direct relevance reasoning), "w/ comparison reasoning" (with comparison reasoning), and "w/ direct relevance & comparison" (with both direct relevance and comparison reasoning) variants.
As shown in Table  1, models without reasoning (using pairwise or listwise loss) perform at a baseline level, with NDCG@5 scores of 73.2% and 73.8 on DL19 and 70.8% and 71.0% on DL20. Adding reasoning significantly boosts performance. For example, direct relevance reasoning improves NDCG@5 to 74.5% on DL19 and 70.1% on DL20 by providing direct relevance relevance explanations. Comparison reasoning, which assesses the relative importance between documents, raises NDCG@5 to 74.1% on DL19 and 72.3% on DL20. Combining both types of reasoning yields the best results, with NDCG@5 scores of 75.4% on DL19 and 72.4% on DL20.
Interestingly, while reasoning improves performance in most cases, there are instances where models without reasoning (e.g., listwise loss) outperform those with reasoning.
4.2.3 Comparison with Baselines
Our student models, trained with direct relevance and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models. As shown in Table  1,the improvements over the baseline models are statistically significant, with p-values less than 0.05.
@@ 4.2.3 Comparison with Baselines
4.2.4 R2R Student Model vs. Baselines
Our R2R student model indicates strong performance across various datasets, often outperforming the baseline models. For instance, on the DL19 dataset, our student model achieves an NDCG@5 of 75.3%, surpassing DeBERTa (68.5%), RankVicuna (71.1%), and RankZephyr (72.2%). Similarly, on the DL20 dataset, our model attains 72.3%, outperforming DeBERTa (64.1%) and RankVicuna (68.6%).
On some datasets, our model achieves the best performance among all models. For example, on the News dataset, our student model reaches an NDCG@5 of 53.8%, slightly outperforming RankZephyr (53.3%) and significantly better than DeBERTa (50.2%).
However, there are cases where other models perform better. On the Touche dataset, RankVicuna achieves an NDCG@5 of 48.7%, which is higher than our student model’s 36.2%. This may be due to RankVicuna’s design for zero-shot listwise reranking, which could be particularly effective for argument retrieval tasks represent in Touche. Similarly, on the Robust04 dataset, RankZephyr attains a higher NDCG@5 of 60.7% compared to our model’s 53.5%.
@@ 4.2.5 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
The BRIGHT dataset Su et al. (2024) is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance. We report the Inst-LSu et al. (2024, 2022) model as a baseline because it is fine-tuned with task-specific instructions, making it particularly well-suited for tasks requiring deeper reasoning. Table 2 shows the NDCG@5 results on the BRIGHT dataset. In the Earth Science domain, our R2R model achieves an NDCG@5 of 31.0%, closely approaching the teacher model’s 37.1%. Similarly, in Psychology, the student model attains 28.9%, compared to the teacher’s 34.2% and BM25’s 11.6%. However, we observe that in certain domains, such as Pony and AoPS, the performance of our student and teacher models is relatively low.
Domain BM25 Inst-L Su et al. (2024) R2R GPT-4 Teacher
Bio. 17.2 21.0 24.3 28.0
Earth. 24.2 27.0 31.0 37.1
Econ. 13.2 16.5 19.0 22.1
Psy. 11.6 26.0 28.9 34.2
Rob. 12.0 14.5 16.0 19.0
Stack 16.0 18.5 21.0 24.2
Sus. 12.4 16.2 18.0 21.9
Leet. 22.1 25.0 27.6 31.4
Pony 8.5 5.2 5.9 6.7
AoPS 6.0 6.8 7.6 9.2
TheoQ. 9.4 15.3 18.5 23.0
TheoT. 3.5 14.2 17.0 21.7
Average 13.1 17.5 19.6 23.2
@@
5 Ablation Study
5.1 Quality of Generated Reasons
@@
Dataset Student Model BLEU ROUGE-L
DL19 w/o Rank Loss 18.0 32.5
w/ Rank Loss 21.4 36.8
DL20 w/o Rank Loss 19.5 34.0
w/ Rank Loss 24.0 38.2
Covid w/o Rank Loss 16.0 30.0
w/ Rank Loss 18.9 34.5
Touche w/o Rank Loss 17.5 33.0
w/ Rank Loss 20.5 37.1
News w/o Rank Loss 20.0 35.0
w/ Rank Loss 22.7 39.3
Table 3: Evaluation of reason generation using BLEU and ROUGE-L scores for student models trained with and without the ranking loss function. The ’Student Model w/o Rank Loss’ is trained without the rank loss function and has the same input as the ’Student Model w/ Rank Loss’.
5.2 Impact of Reasoning Strategies
In Table 4, by removing comparison reasoning results in a performance drop on DL19 from 75.3 to 74.1 (NDCG@5). Similarly, removing direct relevance reasoning causes a performance decrease on DL20 from 72.3 to 72.2 (NDCG@5). Both reasoning strategies contribute to performance improvements.
@@ 5.2 Impact of Reasoning Strategies
Direct relevance & comparison 75.3 / 73.8 72.3 / 71.22
w/ direct relevance reasoning 74.5 / 73.1 70.1 / 70.9
w/ comparison reasoning 74.1 / 72.9 72.2 / 70.4
w/o reasoning 73.2 / 72.1 70.8 / 69.8
Table 4: Ablation study: Impact of direct relevance and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
5.3 Effects of Training Data Size
Figure 4 i shows the performance of small models trained on 100, 1,000, and 2,000 examples. Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it. We find similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
@@ 5.3 Effects of Training Data Size
Models DL19 DL20 BEIR-6 Avg. Covid Touche News NFCorpus Robust04 DBPedia
BM25 50.6 48.0 41.0 59.5 44.2 39.5 30.8 40.7 31.8
Rank GPT-4 Model
Rank GPT-4 (w/o reasoning) 75.6 70.6 54.0 85.5 38.5 57.6 36.6 57.6 47.1
+ pairwise reasoning 75.3 69.9 52.9 83.3 38.9 55.9 36.1 52.9 47.0
+ comparison reasoning 76.3 71.1 53.4 84.6 38.2 56.5 36.5 56.5 48.2
+ direct relevance & comparison 77.7 73.2 54.4 85.3 38.3 58.4 36.3 58.6 49.5
Claude Model
Claude (w/o reasoning) 72.0 68.8 50.8 82.8 36.1 54.4 36.4 50.8 42.5
+ pairwise reasoning 70.6 68.2 50.3 82.2 35.9 55.4 35.6 52.4 42.5
+ comparison reasoning 71.3 69.8 51.1 83.3 36.4 55.1 36.7 51.5 43.2
+ direct relevance & comparison 72.1 70.0 52.0 84.0 37.3 55.5 36.4 52.7 44.9
Gemini Model
Gemini (w/o reasoning) 71.1 68.5 50.1 83.0 35.6 53.7 35.5 50.1 42.9
+ pairwise reasoning 69.1 67.0 51.1 81.6 34.8 52.9 36.9 51.3 43.9
+ comparison reasoning 70.2 68.2 50.2 82.4 35.9 53.4 35.4 51.1 42.9
+ direct relevance & comparison 71.4 68.6 51.0 83.5 37.0 53.8 36.1 51.8 43.7
@@ 6 Discussion
Our experiments show that the R2R framework effectively enhances document reranking by incorporating direct relevance and comparative reasoning. The student models achieve competitive results while significantly reducing computational resources compared to large teacher models. The statistically significant improvements confirm that reasoning strategies effectively enhance model performance. An important aspect of R2R is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings. This dual capability suggests that the model could be applied to other tasks requiring both ranking and reasoning generation.
While our current work centers on document reranking in information retrieval, the integration of reasoning with ranking decisions holds promise for broader applications. For instance, in recommendation systems, the model could rank products or services while providing explanations for its choices, enhancing user trust. Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could improve understandability Jiang et al. (2024). However, future work is needed to explore the model’s performance in these varied contexts and verify its generalizability beyond document reranking McElfresh et al. (2022); Rahmani et al. (2022).
@@ 7 Conclusion
In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates direct relevance and comparative reasoning. Our method leverages the power of large language models to generate interpretable reasons, which are then distilled into smaller, more efficient student models. Our experiments show the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
@@ Appendix D Model Behavior Analysis
In the reasoning generation tasks, the model’s performance varied across direct relevance reasoning, comparison reasoning, and their combination. Below, we provide an analysis of unexpected behavior based on 2,000 queries.
@@
The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we find:
  • Pairwise reasoning: 63 duplicate passage identifiers (3%).
@@
Missing documents were most frequent in direct relevance reasoning:
  • Pairwise reasoning: 642 missing documents (32% of queries).
@@ Appendix F Prompts for Reasoning
Table 10: Basic prompt, direct relevance reasoning prompt, and comparison reasoning prompt.
@@ 3. If there is no direct evidence, try to infer from existing evidence and answer one question:
Direct relevance Reasoning Prompt:
Then, read every passage one by one and find the sentence where the method is the direct answer